OpenStudy (anonymous):

Write an equation of a circle with the given center and radius. center (3, 4) and radius 6

OpenStudy (campbell_st):

the general equation is \[(x - h)^2 + (y - k)^2 = r^2\] which has a centre at (h, k) and radius r so substitute the values you have for the equation

OpenStudy (anonymous):

\[(x-3)^{2}+(y-4)^{2}=6^{2}\]

OpenStudy (anonymous):

simplify dat 6

OpenStudy (anonymous):

36

OpenStudy (anonymous):

yeh

OpenStudy (campbell_st):

that's correct